Find the length and width or a rectangular lot with a perimeter of 88 meters if the length is 4 meters more than the width.
Answer by mducky2(62)

The perimeter of a rectangle is just the sum of the length of its sides. Since opposite sides are the same length, the perimeter is just the sum of double the length plus double the width:
perimeter of a rectangle = 2w + 2l

Since you can only solve an equation fully when there is only one variable, let's represent the length in terms of the width.
w = width
l = length = w + 4

The perimeter becomes:
2w + 2l = 2w + 2(w+4)
= 2w + 2w + 8
= 4w + 8

We know the perimeter equals 88, so we can equate the two.
4w + 8 = 88
w + 2 = 22
w = 20
l = w + 4 = 24

Therefore, the rectangle has a width of 20 and a length of 24.